# Higher-order interactions of multi-layer prompt

### v2

@@ Abstract.
The ”pre-train, prompt” paradigm has successfully evolved in representation learning. While current prompt-tuning methods often introduce learnable prompts, they predominantly treat prompts as isolated, independent components across different network layers. This overlooks the complex and synergistic higher-order interactions that exist between prompts at various hierarchical depths, consequently limiting the expressive power and semantic richness of the prompted model. To address this fundamental gap, we propose a novel framework that explicitly models the Higher-order Interactions of Multi-layer Prompt. Our approach conceptualizes prompts from different layers not as separate entities, but as a cohesive system where their inter-relationships are critical. We design an innovative interaction module that captures these sophisticated, non-linear correlations among multi-layer prompts, effectively modeling their cooperative effects. This allows the model to dynamically aggregate and refine prompt information across the network’s depth, leading to a more integrated and powerful prompting strategy. Extensive experiments on eight benchmark datasets demonstrate that our method, by leveraging these higher-order interactions, consistently surpasses state-of-the-art prompt-tuning baselines. The performance advantage is particularly pronounced in few-shot scenarios, validating that capturing the intricate interplay between multi-layer prompts is key to unlocking more robust and generalizable representation learning.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Web miningccs: Computing methodologies Unsupervised learning
@@ 1. Introduction
To address this issue, recent work introduces prompt tuning as an alternative to fine-tuning. Prompt tuning is initially proposed in NLP and achieving remarkable success (Liu et al., 2023b), and can effectively bridge the gap between pre-training objectives and downstream tasks by incorporating lightweight learnable prompt parameters to modify inputs or outputs, without updating the parameters of the pre-trained model. Inspired by this, pre-training and prompt tuning have been extended to the graph domain (Sun et al., 2023a; Fang et al., 2023; Sun et al., 2022). Compared to fine-tuning, graph prompt exhibits lower computational overhead and better generalization.
Existing graph prompt learning methods can be categorized into two categories based on dependency on pre-training strategies. Pre-training-dependent methods (Sun et al., 2023a; Liu et al., 2023a; Chen et al., 2025) design unified templates for pre-training objectives and downstream tasks to reduce their discrepancy. Pre-training-agnostic methods (Fang et al., 2023; Lee et al., 2024; Yu et al., 2025; Fu et al., 2025) enhance generality by inserting learnable feature prompt vectors into inputs or outputs, requiring no specific pre-training strategy. Further, we classify pre-training-agnostic methods by their prompting mechanisms: (1) Single-step prompt tuning (Fang et al., 2023; Fu et al., 2025; Lee et al., 2024; Yu et al., 2024) employ a direct generation approach where the complete prompt for downstream tasks is produced in a single forward pass. (2) Multi-step prompt tuning (Yu et al., 2025) progressively derives final prompts through iterative inference, mimicking the Chain-of-Thought (CoT) (Wei et al., 2022) rationale to achieve stronger expressiveness.
@@ 2.1. Graph Pre-training
In recent years, graph pre-training techniques have attracted extensive research due to their ability to operate without labeled data (Lu et al., 2021; Zhang and Chen, 2018). These methods are primarily categorized into contrastive learning (Veličković et al., 2018; You et al., 2020) and generative learning (Hou et al., 2022; Zheng et al., 2025). Contrastive learning methods construct multiple views by sampling positive and negative samples, then maximizing the consistency between positive samples. Generative learning methods pre-train encoders by reconstructing node features or graph structures. These approaches employ different pre-training objectives to transfer pre-trained knowledge to various downstream tasks. However, due to the significant gap between downstream tasks and pre-training objectives (Sun et al., 2023a, b), the performance on downstream tasks may be compromised.
@@ 2.2. Graph Prompt-Tuning
Recent advances in graph prompt learning have sought to bridge the gap between pre-training and downstream tasks by introducing task-specific prompts. Early works, such as GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a), reformulate downstream tasks as link prediction or subgraph similarity tasks. All-in-One (Sun et al., 2023a) unified various downstream tasks as graph-level tasks. Feature-based prompt method GPF+ (Fang et al., 2023) and SUPT (Lee et al., 2024) introduced feature-space prompting and subgraph prompt, where learnable vectors modify input. EdgePrompt (Fu et al., 2025) inserts edge-level prompts in the aggregation at each level from an edge perspective. ProNOG (Yu et al., 2024) and DAGPrompT (Chen et al., 2025) accounted for node heterophily in the prompt by preserving neighbourhood similarity. The aforementioned methods all belong to single-step prompt tuning approaches. GCOT (Yu et al., 2025) introduces the chain of thought to generate multi-step prompts. However, these methods only focus on single-granularity information during prompt generation, ignoring the rich multi-granularity features that co-exist in real-world graphs.
@@ 4.2. Multi-scale Prompt
where CONDNET is a conditional network parameterised by the parameter $\phi$. The conditional network can be viewed as a hypernetwork (Zhou et al., 2022), i.e., a lightweight auxiliary network. To enhance task-relevant features while suppressing irrelevant noise in the pre-trained representations. Then we use the prompt to modify the original features and obtain prompted embeddings via a pre-trained encoder:
@@ 4.3. Prompt Tuning
In the prompt tuning phase, we follow the formulation of  (Liu et al., 2023a), which defines the task within a subgraph similarity framework. Accordingly, our loss function remains consistent with this design and is specified as follows:
@@ 5.1. Experinmental Setting
Datasets. We evaluate the performance of MSGCOT using several datasets from different domains, including citation networks, e-commerce, protein structures, and molecular graphs. For the node classification task, we used three citation network datasets, Cora (McCallum et al., 2000), Citeseer, Pubmed (Sen et al., 2008), and the e-commerce co-purchase network Photo (Shchur et al., 2018). For graph classification, we used a protein structure dataset PROTEINS (Borgwardt et al., 2005), a molecular graph dataset MUTAG, BZR, and COX2 (Rossi and Ahmed, 2015), which represent, respectively, nitro-aromatic compounds, ligands related to benzodiazepine receptors, and molecular structures related to cyclo-oxidase-2 inhibitors.
@@
Baselines. We compare MSGCOT with current state-of-the-art methods. These baselines fall into three categories:(1) Supervised learning. We use a supervised approach on GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017) to train on downstream labels. (2) Pre-training + fine-tuning. LP (Zhang and Chen, 2018), GraphCL (You et al., 2020), and DGI/InfoGraph (Veličković et al., 2018; Sun et al., 2019) employ a pre-training and fine-tuning strategy, where the model is pre-trained on unlabeled data and then fine-tuned on downstream tasks. (3) Pre-training + Prompt: We uniformly use the link prediction task to pre-train the GNNs, and use graph prompt methods for prompting, including single-step prompt methods All-in-One (Sun et al., 2023a), GPF, GPF+ (Fang et al., 2023), SUPT (Lee et al., 2024), GraphPrompt (Liu et al., 2023a), EdgePrompt, EdgePrompt+ (Fu et al., 2025), DAGPrompT (Chen et al., 2025), and multi-step prompt method GCOT (Yu et al., 2025).
@@
Implementation details. We evaluate our approach on few-shot node classification and graph classification tasks. For each class, we randomly select $k$ nodes or graphs as the training set. For node classification, following the setup of GCOT, we randomly sample 1,000 nodes from the remaining nodes as the test set. For graph classification, we use the remaining samples as the test set. We report the accuracy of the models. To ensure a fair comparison, all prompt-based models adopt link prediction as the pre-training strategy, and GCN is used as the backbone with a hidden layer size of 256. The number of coarsening layers is set to 2, and the coarsening ratio is selected from {0.01, 0.1, 0.2, 0.3}. For the hidden dimension of the low-rank decomposition matrix, we set it to 8. For the prompt constraint weight $\alpha$, we set it to 1 in the node classification and 0 in the graph classification. We conduct 100 random sampling trials, generating 1 k-shot tasks for both node classification and graph classification, and report the mean and variance of the results. The pre-training strategies and hyperparameter settings of other baseline methods are based on their original papers. All experiments are conducted on an NVIDIA RTX 4090D GPU with 24GB of memory. We also provide experimental results using graphCL as a pre-training task in Appendix 2.2. See Appendix 1.2 for detailed parameter settings. For code implementation, consult the supplementary materials.
@@ 5.4. Ablation study
As shown in the Table 2, (1) Multi-scale prompts are crucial for performance improvement. After removing it, the accuracy of node classification tasks decreased by an average of 5.52%. In contrast, graph classification tasks experienced a more pronounced performance drop, averaging 17.7%, validating the critical role of hierarchical prompts in capturing multi-scale structural features. (2) The reconstruction loss significantly impacts node classification tasks, while graph classification focuses more on global information. (3) The backtracking mechanism is particularly critical for graph classification, with performance dropping sharply by 12-15% when using unidirectional prompts, confirming the importance of dynamically adjusting prompt granularity; (4) Progressive prompt updates yield a performance gain of 2-5%, indicating that intermediate-layer prompts carry complementary hierarchical information. These findings demonstrate that MSGCOT can coordinate features at different granularities, verifying the effectiveness of a multi-level, multi-scale prompt.
@@ 5.5. Hyperparameter Analysis
Coarsening rate. We employ the coarsening rate to control the number of coarsened nodes, where a smaller coarsening rate indicates fewer coarsened nodes and places greater emphasis on coarse-grained information. We validate the impact of different coarsening rates in Figure 4. Our findings reveal that the two tasks exhibit significantly distinct patterns: (1) For node-level tasks, the best performance occurs at moderate coarsening rates (0.1-0.3). When the coarsening rate is too small, critical node information is lost, leading to performance degradation. (2) For graph-level tasks, graph classification maintains stable high performance across a wider range of coarsening rates (0.05-0.3), demonstrating stronger robustness to coarsening.
@@
Coarsening layer. The number of coarsening layers represents both the count of multi-scale prompt layers and the steps prompt. At each prompting step, hierarchical prompts with different granularities are injected. We validate the impact of varying layer numbers on multi-scale prompting. As shown in Figure 5, we can have the following observations. (1) For node-level tasks, performance first increases and then decreases with additional layers. On the Cora dataset, optimal performance consistently appears at 2 layers, demonstrating that this configuration achieves the best balance between node-level semantics and multi-layer subgraph semantics. Excessive layers may cause information loss due to over-smoothing. (2) For graph-level tasks, performance shows continuous improvement in the initial stages as layers increase. On the PROTEINS dataset, when L=3 or 4, the model still maintains strong performance, indicating that deeper architectures can provide additional benefits. This reflects graph-level tasks’ stronger dependence on hierarchical features and validates the effectiveness of our hierarchical prompt injection mechanism.
@@
[…]
@@ 5.7. Parameter and Efficiency Analysis
Trainable prompt parameters.We compare the trainable parameter counts of different prompting methods across various datasets. The experimental results in Table 4 demonstrate that MSGCOT significantly reduces trainable parameters while maintaining model performance: For node classification tasks, its parameter count (10.37K-21.8K) represents a 47.1%-68.3% reduction compared to GCOT (32.76K), while being comparable to EdgePrompt+. The improvement is even more pronounced in graph classification tasks, where the parameter count (1.17K-5.81K) shows a 29.1%-85.7% reduction versus GCOT (8.19K), even outperforming single-step methods like GPF+ (5.12K) on datasets such as MUTAG. This remarkable parameter efficiency stems from MSGCOT’s innovative low-rank design (r=8 for node classification, r=1 for graph classification) and task-adaptive hierarchical coarsening mechanism, validating its advantage in balancing model complexity and performance. On average, the parameter reduction reaches 53.6% for node classification and 63.2% for graph classification tasks.
@@
Running times. We compare the per-epoch running time of different prompt tuning methods across multiple datasets. The experimental results in Table 5 demonstrate that MSGCOT exhibits distinct efficiency advantages across different task types. For node classification tasks, MSGCOT’s per-epoch runtime is slightly higher than GPF+ and EdgePrompt+, but remains comparable to GCOT. In graph classification tasks, it achieves an average per-epoch runtime of only 0.137 seconds, a 34.8% reduction compared to GCOT, while maintaining significant accuracy advantages. This stems from either the extremely low coarsening ratio or the reduced hidden dimension in graph classification tasks. Although multi-step prompting inevitably introduces additional complexity, its runtime overhead remains within a controllable range.
@@ 6. Conclusion
In this paper, we propose MSGCOT, the first novel multi-scale prompt chain framework, thus addressing the key limitation of prompt granularity singularity in existing graph prompt tuning. The proposed low-rank coarsening network effectively captures hierarchical structural features while maintaining parameter efficiency. The method mimics the human thought process by gradually introducing different information from coarse to fine granularity during the prompt generation, ultimately generating more refined prompts. Experimental results show that compared to methods that only consider a single granularity, MSGCOT exhibits better performance, significantly outperforming existing benchmark methods.

### v3

# Beyond Single-Granularity Prompts: A Multi-Scale Chain-of-Thought Prompt Learning for Graph

@@ Abstract.
The “pre-train, prompt” paradigm, designed to bridge the gap between pre-training tasks and downstream objectives, has been extended from the NLP domain to the graph domain and has achieved remarkable progress. Current mainstream graph prompt-tuning methods modify input or output features using learnable prompt vectors. However, existing approaches are confined to single-granularity (e.g., node-level or subgraph-level) during prompt generation, overlooking the inherently multi-scale structural information in graph data, which limits the diversity of prompt semantics. To address this issue, we pioneer the integration of multi-scale information into graph prompt and propose a Multi-Scale Graph Chain-of-Thought (MSGCOT) prompting framework. Specifically, we design a lightweight, low-rank coarsening network to efficiently capture multi-scale structural features as hierarchical basis vectors for prompt generation. Subsequently, mimicking human cognition from coarse-to-fine granularity, we dynamically integrate multi-scale information at each reasoning step, forming a progressive coarse-to-fine prompt chain. Extensive experiments on eight benchmark datasets demonstrate that MSGCOT outperforms the state-of-the-art single-granularity graph prompt-tuning method, particularly in few-shot scenarios, showcasing superior performance. The code is available at: https://github.com/zhengziyu77/MSGCOT.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792115isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Web miningccs: Computing methodologies Unsupervised learning
@@ 1. Introduction
To address this issue, recent work introduces prompt tuning as an alternative to fine-tuning. Prompt tuning is initially proposed in NLP and achieving remarkable success (Liu et al., 2023a), and can effectively bridge the gap between pre-training objectives and downstream tasks by incorporating lightweight learnable prompt parameters to modify inputs or outputs, without updating the parameters of the pre-trained model. Inspired by this, pre-training and prompt tuning have been extended to the graph domain (Sun et al., 2023a; Fang et al., 2023; Sun et al., 2022). Compared to fine-tuning, graph prompt exhibits lower computational overhead and better generalization.
Existing graph prompt learning methods can be categorized into two categories based on dependency on pre-training strategies. Pre-training-dependent methods (Sun et al., 2023a; Liu et al., 2023b; Chen et al., 2025) design unified templates for pre-training objectives and downstream tasks to reduce their discrepancy. Pre-training-agnostic methods (Fang et al., 2023; Lee et al., 2024; Yu et al., 2025; Fu et al., 2025) enhance generality by inserting learnable feature prompt vectors into inputs or outputs, requiring no specific pre-training strategy. Further, we classify pre-training-agnostic methods by their prompting mechanisms: (1) Single-step prompt tuning (Fang et al., 2023; Fu et al., 2025; Lee et al., 2024; Yu et al., 2024) employ a direct generation approach where the complete prompt for downstream tasks is produced in a single forward pass. (2) Multi-step prompt tuning (Yu et al., 2025) progressively derives final prompts through iterative inference, mimicking the Chain-of-Thought (CoT) (Wei et al., 2022) rationale to achieve stronger expressiveness.
@@ 2.1. Graph Pre-training
In recent years, graph pre-training have attracted extensive research due to their ability to operate without labeled data (Lu et al., 2021; Zhang and Chen, 2018). These methods are primarily categorized into contrastive learning (Veličković et al., 2018; You et al., 2020) and generative learning (Hou et al., 2022; Zheng et al., 2025). Contrastive learning methods construct multiple views by sampling positive and negative samples, then maximizing the consistency between positive samples. Generative learning methods pre-train encoders by reconstructing node features or graph structures. These approaches employ different pre-training objectives to transfer pre-trained knowledge to various downstream tasks. However, due to the significant gap between downstream tasks and pre-training objectives (Sun et al., 2023a, b), the performance on downstream tasks may be compromised.
@@ 2.2. Graph Prompt-Tuning
Recent advances in graph prompt learning have sought to bridge the gap between pre-training and downstream tasks by introducing task-specific prompts. Early works, such as GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023b), reformulate downstream tasks as link prediction or subgraph similarity tasks. All-in-One (Sun et al., 2023a) unified various downstream tasks as graph-level tasks. Feature-based prompt method GPF+ (Fang et al., 2023) and SUPT (Lee et al., 2024) introduced feature-space prompting and subgraph prompt, where learnable vectors modify input. EdgePrompt (Fu et al., 2025) inserts edge-level prompts in the aggregation at each level from an edge perspective. ProNOG (Yu et al., 2024) and DAGPrompT (Chen et al., 2025) accounted for node heterophily in the prompt by preserving neighbourhood similarity. The aforementioned methods all belong to single-step prompt tuning approaches. GCOT (Yu et al., 2025) introduces the chain of thought to generate multi-step prompts. However, these methods only focus on single-granularity information during prompt generation, ignoring the rich multi-granularity features that co-exist in real-world graphs.
@@ 4.2. Multi-scale Prompt
where CONDNET is a conditional network by the parameter $\phi$. The conditional network can be viewed as a hypernetwork (Zhou et al., 2022), i.e., a lightweight auxiliary network. To enhance task-relevant features while suppressing irrelevant noise in the pre-trained representations. Then we use the prompt to modify the original features and obtain prompted embeddings via a pre-trained encoder:
@@ 4.3. Prompt Tuning
In the prompt tuning phase, we follow the formulation of  (Liu et al., 2023b), which defines the task within a subgraph similarity framework. Accordingly, our loss function remains consistent with this design and is specified as follows:
@@ 5.1. Experinmental Setting
Datasets. We evaluate the performance of MSGCOT using several datasets from different domains, including citation networks, e-commerce, protein structures, and molecular graphs. For the node classification task, we used three citation network datasets, Cora (McCallum et al., 2000), Citeseer, Pubmed (Sen et al., 2008), and the e-commerce co-purchase network Photo (Shchur et al., 2018). For graph classification, we used a protein structure dataset PROTEINS (Borgwardt et al., 2005), a molecular graph dataset MUTAG, BZR, and COX2 (Rossi and Ahmed, 2015).
@@
Baselines. We compare MSGCOT with current state-of-the-art methods. These baselines fall into three categories:(1) Supervised learning. We use a supervised approach on GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017) to train on downstream labels. (2) Pre-training + fine-tuning. LP (Zhang and Chen, 2018), GraphCL (You et al., 2020), and DGI/InfoGraph (Veličković et al., 2018; Sun et al., 2019) employ a pre-training and fine-tuning strategy, where the model is pre-trained on unlabeled data and then fine-tuned on downstream tasks. (3) Pre-training + Prompt: We uniformly use the link prediction task to pre-train the GNNs, and use graph prompt methods for prompting, including single-step prompt methods All-in-One (Sun et al., 2023a), GPF, GPF+ (Fang et al., 2023), SUPT (Lee et al., 2024), GraphPrompt (Liu et al., 2023b), EdgePrompt, EdgePrompt+ (Fu et al., 2025), DAGPrompT (Chen et al., 2025), and multi-step prompt method GCOT (Yu et al., 2025).
@@
Implementation details. We evaluate our approach on a few-shot node classification and graph classification tasks. For each class, we randomly select $k$ nodes or graphs as the training set. For node classification, following the GCOT setup, we randomly sample 1,000 nodes from the remaining nodes as the test set. For graph classification, we use the remaining samples as the test set. We report the model accuracy. To ensure a fair comparison, all prompt-based models adopt link prediction as the pre-training strategy, using GCN as the backbone with a hidden layer size of 256. The number of coarsening layers is set to 2, and the coarsening ratio is selected from 0.01,0.1,0.2,0.3. The hidden dimension of the low-rank decomposition matrix is set to 8. The prompt constraint weight $\alpha$ is set to 1 for node classification and 0 for graph classification. We conduct 100 random sampling trials, generating one-shot tasks for both node classification and graph classification, and report the mean and variance of the results. The pre-training strategies and hyperparameter settings for other baseline methods follow their original papers. All experiments are conducted on an NVIDIA RTX 4090D GPU with 24GB of memory. Experimental results using GraphCL as a pre-training task are provided in Appendix B.1. Detailed parameter settings can be found in Appendix A.2.
@@ 5.4. Ablation study
As shown in Table 2: (1) Multi-scale prompts are crucial for performance. Their removal reduces node classification accuracy by 5.52% on average and leads to a more pronounced drop in graph classification (17.7% on average), validating their critical role in capturing hierarchical structural features. (2) The reconstruction loss significantly impacts node classification, while graph classification relies more on global information. (3) The backtracking mechanism is particularly vital for graph classification; performance drops sharply by 12-15% when using unidirectional prompts, confirming the importance of dynamically adjusting prompt granularity. (4) Progressive prompt updates yield a 2-5% performance gain, indicating that intermediate-layer prompts carry complementary hierarchical information. These results demonstrate that MSGCOT effectively coordinates features across different granularities, verifying the efficacy of its multi-level, multi-scale prompt design.
@@ 5.5. Hyperparameter Analysis
Coarsening rate. We control the number of coarsened nodes via the coarsening rate, where lower values emphasize coarse-grained information. As shown in Figure 4, node-level tasks peak at moderate rates (0.1–0.3), as excessively low rates lose fine-grained details. Graph classification remains robust across a wider range (0.05–0.3), demonstrating stronger robustness to coarsening.
@@
Coarsening layer. The number of coarsening layers defines both the depth of multi-scale prompting and the reasoning steps. At each step, hierarchical prompts of varying granularity are integrated. We analyze the effect of layer count in Figure 5 and observe: (1) For node-level tasks, performance initially improves then declines with more layers. On Cora, two layers achieve an optimal balance between node semantics and subgraph structure; excessive layers induce over-smoothing. (2) For graph-level tasks, performance steadily improves with deeper hierarchies. On PROTEINS, performance remains strong at 3–4 layers, indicating enhanced benefits from hierarchical feature extraction. This confirms graph tasks’ greater reliance on multi-scale structures and validates our hierarchical prompting mechanism.
@@
Hidden dimensionality of coarsening net. We design an efficient coarsening network to capture multi-scale information and improve generalization by controlling parameter scale. To assess its efficacy, we evaluate performance across different hidden dimensions $r$ (Figure 6). For node classification, accuracy stabilizes when $r>8$, with peak results at $r=8$. Graph classification exhibits greater robustness to dimension changes, achieving strong performance even at $r=1$. A slight overfitting emerges in node tasks when $r>32$. These results confirm that MSGCOT’s low-rank design effectively balances efficiency and performance stability, sustaining strong generalization with minimal dimensionality.
@@ 5.7. Parameter and Efficiency Analysis
Trainable prompt parameters.We compare trainable parameters across prompting methods in Table 4. MSGCOT maintains competitive performance while substantially reducing parameters: for node classification (10.37K–21.8K), it achieves a 47.1%–68.3% reduction compared to GCOT (32.76K). In graph classification (1.17K–5.81K), reductions reach 29.1%–85.7% versus GCOT (8.19K), even outperforming GPF+ (5.12K) on MUTAG. This efficiency stems from MSGCOT’s low-rank design (r=8 for node tasks, r=1 for graph tasks) and task-adaptive coarsening. On average, parameter reductions are 53.6% (node) and 63.2% (graph), highlighting its ability to balance complexity and performance.
@@
Running times. Table 5 compares the per-epoch running times of different prompt tuning methods. MSGCOT shows clear efficiency benefits across task types. For node classification, its runtime is slightly higher than GPF+ and EdgePrompt+ but comparable to GCOT. For graph classification, MSGCOT averages 0.137 seconds per epoch—a 34.8% reduction over GCOT—while maintaining higher accuracy, attributable to its low coarsening ratio and reduced hidden dimension. Although multi-step prompting adds complexity, MSGCOT’s overhead remains manageable.
@@ 6. Conclusion
In this paper, we propose MSGCOT, a novel multi-scale prompt chain framework that addresses the limitation of single-granularity prompts in existing graph prompt tuning methods. The proposed low-rank coarsening network captures hierarchical structural features while remaining parameter-efficient. Inspired by the human thought process, MSGCOT progressively introduces information from coarse to fine granularity during prompt generation, leading to more refined prompts. Experimental results demonstrate that MSGCOT consistently outperforms single-granularity approaches and achieves superior performance.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants 62425605, 62133012, and 62303366, and in part by the Key Research and Development Program of Shaanxi under Grants 2025CY-YBXM-041, 2022ZDLGY01-10, and 2024CY2-GJHX-15.